# Monadic reconstruction of unitary Drinfeld centers and Factorization Homology

### v1

@@ 4.2. Unitary modules
$\mathbb{A}(a_{1})\ni f\mapsto\alpha\circ(f\odot\mathrm{id}_{H})\in\operatorname{Vec}(\mathcal{C}(a\odot H,H))$
@@
This definition of unitary modules was already used in [MR3687214], Section 5, but we make it explicit how it relates to the standard notion of modules for algebra objecst.
@@
The previous Lemma says that $\operatorname{LMod}_{\mathrm{Hilb}(\mathcal{C})}(\mathbb{A})$ is a right $\mathrm{Hilb}(\mathcal{C})$-module category, and it is straightforward to check that it has a canonical structures of right $\mathrm{Hilb}(\mathcal{C})$-module W-category.
@@ 5. The Canonical $W^{*}$-Algebra
Every unitary tensor category $\mathcal{C}$ is canonically a right $(\mathcal{C}^{mp}\boxtimes\mathcal{C})$-module C-category, with the action is given by
@@
is an equivalence of right $(\mathcal{C}^{mp}\boxtimes\mathcal{C})$-module C-categories, where $\mathcal{M}_{\mathcal{S}}$ is the Cauchy completion of the category of free $\mathcal{S}$-left modules of the form $\mathcal{S}\odot(a\boxtimes b)$, with $a,b\in\mathcal{C}$.
@@
From now on, $\mathcal{C}$ is a fixed unitary tensor category and $\mathcal{S}$ is its canonical W-algebra.
@@
It remains to show that every left $\mathcal{S}$-module in $\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})$ decomposes as a direct sum as above. Recall the adjunction between the free $\mathcal{S}$-module functor and the forgetful functor from $\operatorname{LMod}(\mathcal{S})\to\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})$: given a left $\mathcal{S}$-module $V$ in $\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})$,
$\operatorname{Hom}_{\operatorname{LMod}(\mathcal{S})}(\mathcal{S}\odot(b\boxtimes c),V)\simeq\operatorname{Hom}_{\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})}(b\boxtimes c,V)\ .$
@@
$\mathcal{S}_{H}(b\boxtimes c)\simeq\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}\mathcal{S}_{a}(b\boxtimes c)\odot H(a)\simeq\bigoplus_{a\leq bc}\mathcal{C}(bc,a)\odot H(a)\simeq H(bc)\ .$
@@
$\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})((\overline{a}\boxtimes a)\odot\mathcal{S}_{\mathds{1}},\mathcal{S}_{\mathds{1}})\simeq\prod_{a_{1},a_{2}}\mathcal{L}\left(\mathcal{C}(\overline{a}_{1},\overline{a}_{2}\overline{a})\odot\mathcal{C}(a_{1},aa_{2})\odot\mathcal{S}(\bar{a}_{2}\boxtimes a_{2}),\mathcal{S}(\overline{a_{1}}\boxtimes a_{1})\right)\ .$
@@
Observe now that, as algebra objects, $\mathcal{L}^{l}_{\mathcal{S}_{\mathds{1}}}=\mathbb{B}^{l}_{\mathcal{S}_{\mathds{1}}}$.
@@
it follows that $\Gamma_{a_{1},a_{2}}(\iota_{a})\|\geq 1$ but an easy application of the Cauchy-Schwarz inequality shows that $\|\Gamma_{a_{1},a_{2}}(\iota_{a})\|\leq 1$. ∎
@@
$H(a)=\bigoplus_{b,c}^{\ell^{2}}\mathcal{C}(a,bc)\otimes\mathcal{K}(b\boxtimes c)\ ,$
@@ 5.1. Bimodules and half-braidings
Recall that a half-braiding on $H$ can be described as a collection of invertible linear maps $H(c\bar{a})\to H(\bar{a}c)$ indexed by $a,c\in\operatorname{Irr}(\mathcal{C})$, satisfying certain relations. Since $\mathcal{S}(a\boxtimes b)\simeq\delta_{a,\bar{b}}\mathbb{C}$, $\alpha$ is determined by algebra homomorphisms
@@
$\mathcal{S}_{H}\boxtimes\mathcal{S}_{K}=\mathcal{S}_{H\otimes K}\ .$
@@ 6.1. Reduction to centrally pointed bimodules
It was proved in [hataishi2025categorical] that centrally pointed $\mathcal{C}$-bimodule C-categories give rise to C-algebra objects in $\mathcal{Z}\operatorname{Vec}(\mathcal{C})$, i.e. C-algebra objects in $\operatorname{Vec}(\mathcal{C})$ equipped with braidings which is compatible with the involutions. We call such C-algebra objects internal Yetter-Drinfled C-algebras. When $\mathrm{Hilb}(\mathcal{C})=\operatorname{Rep}(\mathbb{K})$ for a compact quantum group $\mathbb{K}$, the category of internal Yetter-Drinfeld C-algebras is equivalent to the category of continuous $D\mathbb{K}$-C-algebras. In this case, $\operatorname{Rep}(D\mathbb{K})\simeq\mathcal{Z}\operatorname{Rep}(\mathbb{K})$, as unitarily braided C-tensor categories.
@@ 6.3. Realization of Yetter-Drinfeld C-algebra objects
of C-algebras, called the realization of $\mathbb{D}$ along $F$, or the cross-product of $A$ by $(F,\mathbb{D})$. This construction is functorial in the appropriate sense, see [Hataishi_Palomares_2025].
Every unitary tensor functor $\tilde{F}:\mathcal{C}\to{\sf Bim_{fgp}}(A)$ induces a unitary tensor functor $F:\mathcal{C}^{mp}\boxtimes\mathcal{C}\to{\sf Bim_{fgp}}(A^{\text{op}}\otimes A)$, where $\otimes$ is either the minimal or the maximal tensor product of C-algebras. $F$ is constructued from $\tilde{F}$ as follows. Given $X\in{\sf Bim_{fgp}}(A)$, let $X^{\#}$ be the complex conjugate vector space. Then
@@
Moreover, of $\psi:(\mathcal{M},m)\to(\mathcal{N},n)$ is a morphism of pointed right $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C-categories, then there is an induced *-homomorphism $\psi_{*}:(A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{m}\to(A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{n}$ the fixes $B_{F}$ point-wise. The operation $\psi\mapsto\psi_{*}$ respects composition, and we obtain therefore a functor from the category of pointed right $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C-categories to extensions of the symmetric enveloping algebra $B_{F}$.
@@ 6.4. Realization of Categories of Internal Modules
Let $\mathcal{D}$ be a unitary tensor category, $\mathbb{D}$ a C-algebra object in $\operatorname{Vec}(\mathcal{D})$, and suppose $F:\mathcal{D}\to\operatorname{Corr}(B)$ is a fully-faithful unitary tensor functor. Recall that $\mathcal{M}_{\mathbb{D}}$ denotes the Cauchy completion of the category of free left $\mathbb{D}$-modules of the form $\mathbb{D}\odot a$, with $a\in\mathbb{D}$.
@@
Fiven $\xi_{c}\odot\eta_{c}\in F(c)\odot\mathbb{D}(c)$ and $\xi_{b}\odot\eta_{b}\in F(b)\odot\mathbb{D}_{a}(b)$, define
@@
See [Hataishi_Palomares_2025]. The following Lemma is straightforward.
@@
$\langle\xi_{b_{1}}\odot\eta_{b_{1}},\xi_{b_{2}}\odot\eta_{b_{2}}\rangle:=\sum_{c\in\operatorname{Irr}(\mathcal{D})}\sum_{v\in O(c,b_{1}\bar{b}_{2})}F(v^{*})\left(F_{2}(\xi_{b_{1}}\odot j^{F}(\xi_{b_{2}}))\right)\odot\mathbb{D}(v)\left(\mathbb{D}_{2}((\eta_{b_{1}}\odot j^{\mathbb{D}}(\eta_{b_{2}})(\mathrm{id}_{b_{1}}\odot R_{a}\odot\mathrm{id}_{\bar{b}_{2}})))\right)$
@@
in $B\otimes\mathbb{D}({\mathds{1}})$. There exists a state $\omega$ on $\mathbb{D}({\mathds{1}})$ such that
@@
Then $\Im(\tilde{\iota}_{\omega})\subset\Im(\iota_{\omega})$. The ineqiality $(\mathrm{id}_{B}\otimes\omega)(\mathbb{E}(\langle T(\zeta),T(\zeta)\rangle))>0$ now implies that $\tilde{\iota}_{\omega}(T)\neq 0$, so $T$ is a non-zero operator in $\Im(\iota_{\omega})\simeq\mathcal{M}_{\mathbb{D}}(\mathbb{D}_{a_{1}},\mathbb{D}_{a_{2}})$. ∎

### v2

@@ 4.2. Unitary modules
$\mathbb{A}(a_{1})\ni f\mapsto\alpha\circ(f\odot\mathrm{id}_{H})\in\operatorname{Vec}(\mathcal{C})(a_{1}\odot H,H)$
@@
This definition of unitary modules was already used in [MR3687214], Section 5. Here, we make explain how it relates to the standard notion of modules for algebra objects.
@@
The previous Lemma says that $\operatorname{LMod}_{\mathrm{Hilb}(\mathcal{C})}(\mathbb{A})$ is a right $\mathrm{Hilb}(\mathcal{C})$-module category, and it is straightforward to check that it has a canonical structure of a right $\mathrm{Hilb}(\mathcal{C})$-module W-category.
@@ 5. The Canonical $W^{*}$-Algebra
Every unitary tensor category $\mathcal{C}$ is canonically a right $(\mathcal{C}^{mp}\boxtimes\mathcal{C})$-module C-category, where the action is given by
@@
is an equivalence of right $(\mathcal{C}^{mp}\boxtimes\mathcal{C})$-module W-categories, where $\mathcal{M}_{\mathcal{S}}$ is the Cauchy completion of the category of free $\mathcal{S}$-left modules of the form $\mathcal{S}\odot(a\boxtimes b)$, with $a,b\in\mathcal{C}$.
@@
It remains to show that every left $\mathcal{S}$-module in $\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})$ decomposes as a direct sum as above. Recall the adjunction between the free $\mathcal{S}$-module functor and the forgetful functor from $\operatorname{LMod}_{\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})}(\mathcal{S})\to\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})$: given a left $\mathcal{S}$-module $V$ in $\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})$,
$\operatorname{Hom}_{\operatorname{LMod}_{\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})}(\mathcal{S})}(\mathcal{S}\odot(b\boxtimes c),V)\simeq\operatorname{Hom}_{\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})}(b\boxtimes c,V)\ .$
@@
$\mathcal{S}_{H}(b\boxtimes c)\simeq\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}\mathcal{S}_{a}(b\boxtimes c)\odot H(a)\simeq\bigoplus_{a\leq bc}\mathcal{C}(bc,a)\odot H(a)\simeq H(bc)\ ,$
where the notation $a\leq bc$ means that $a$ is a subobject of $bc$. ∎
@@
$\mathcal{L}^{l}_{\mathcal{S}_{\mathds{1}}}(\bar{a}\boxtimes a)=\operatorname{Vec}(\mathcal{C}^{mp}\boxtimes\mathcal{C})((\overline{a}\boxtimes a)\odot\mathcal{S}_{\mathds{1}},\mathcal{S}_{\mathds{1}})\simeq\prod_{a_{1},a_{2}}\mathcal{L}\left(\mathcal{C}(\overline{a}_{1},\overline{a}_{2}\overline{a})\odot\mathcal{C}(a_{1},aa_{2})\odot\mathcal{S}(\bar{a}_{2}\boxtimes a_{2}),\mathcal{S}(\overline{a_{1}}\boxtimes a_{1})\right)\ .$
@@
it follows that $\Gamma_{a_{1},a_{2}}(\iota_{a})\|\geq 1$ but an application of the Cauchy-Schwarz inequality shows that $\|\Gamma_{a_{1},a_{2}}(\iota_{a})\|\leq 1$. ∎
@@
$H(a)=\bigoplus_{b,c\in\operatorname{Irr}(\mathcal{C})}^{\ell^{2}}\mathcal{C}(a,bc)\otimes\mathcal{K}(b\boxtimes c)\ ,$
@@ 5.1. Bimodules and half-braidings
Since $\mathcal{S}(a\boxtimes b)\simeq\delta_{a,\bar{b}}\mathbb{C}$, $\alpha$ is determined by algebra homomorphisms
@@
Since the right-hand side does not depend on $b$ or $c$, this gives a uniform bound on the norms of the components of $\alpha(\iota_{a})$. ∎
@@
$\mathcal{S}_{H}\otimes\mathcal{S}_{K}=\mathcal{S}_{H\otimes K}\ .$
@@ 6. $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C-categories
Consider the 2-category $\operatorname{C*Cat}$ of C-categories: objects are small C-categories, 1-morphisms are *-functors, and 2-morphisms are unitary natural isomorphisms. This category was studied in detail in [MR3007088], where the following is shown.
Theorem 6.1 ([MR3007088], Section 4).
Remark 6.2.
Small C-tensor categories are exactly the algebras in $\operatorname{C*Cat}$. Given a C-tensor category $\mathcal{D}$, left and right $\mathcal{D}$-module C-categories $\mathcal{M}$ and $\mathcal{N}$, it follows from Dell’Ambrogio’s results that the functor
$\operatorname{C*Cat}\ni\mathcal{E}\mapsto\operatorname{Bal}_{\mathcal{D}}(\mathcal{M}\boxtimes\mathcal{N},\mathcal{E})\ ,$
where $\operatorname{Bal}_{\mathcal{D}}(\mathcal{M}\boxtimes\mathcal{N},\mathcal{E})$ denotes the C-category of $\mathcal{D}$-balanced unitary functors from $\mathcal{M}\boxtimes\mathcal{N}$ to $\mathcal{E}$, is representable: there exists a C-category $\mathcal{M}\underset{\mathcal{D}}{\boxtimes}\mathcal{N}$ such that
$\operatorname{Bal}_{\mathcal{D}}(\mathcal{M}\boxtimes\mathcal{N},\mathcal{E})\simeq\operatorname{C*Cat}(\mathcal{M}\underset{\mathcal{D}}{\boxtimes}\mathcal{N},\mathcal{E})$
as C-categories.
Definition 6.3.
The C-category $\mathcal{M}\underset{\mathcal{D}}{\boxtimes}\mathcal{N}$, defined up to unitary isomorphism, is called the balanced tensor product of $\mathcal{M}$ and $\mathcal{N}$ over $\mathcal{D}$.
@@ 6.1. Reduction to centrally pointed bimodules
It was proved in [hataishi2025categorical] that centrally pointed $\mathcal{C}$-bimodule C-categories give rise to C-algebra objects in $\mathcal{Z}\operatorname{Vec}(\mathcal{C})$, i.e. C-algebra objects in $\operatorname{Vec}(\mathcal{C})$ equipped with braidings which are compatible with the respective involutions. We call such C-algebra objects internal Yetter-Drinfled C-algebras. When $\mathrm{Hilb}(\mathcal{C})=\operatorname{Rep}(\mathbb{K})$ for a compact quantum group $\mathbb{K}$, the category of internal Yetter-Drinfeld C-algebras is equivalent to the category of continuous $D\mathbb{K}$-C-algebras. In this case, $\operatorname{Rep}(D\mathbb{K})\simeq\mathcal{Z}\operatorname{Rep}(\mathbb{K})$, as unitarily braided C-tensor categories.
@@ 6.3. Realization of Yetter-Drinfeld C-algebra objects
Given a C-algebra AA, we write ${\sf Bim_{fgp}}(A)$ for the category of finitely generated projective Hilbert C-bimodules over $A$. See [Hataishi_Palomares_2025] for more details.
@@
of C-algebras, called the realization of $\mathbb{D}$ along $F$, or the cross-product of $A$ by $(F,\mathbb{D})$. It comes with a canonical faithful conditional expectation $\mathbb{E}:A\rtimes\mathbb{D}\to A\otimes\mathbb{D}({\mathds{1}})$. This construction is functorial in the appropriate sense, see [Hataishi_Palomares_2025].
Every unitary tensor functor $\tilde{F}:\mathcal{C}\to{\sf Bim_{fgp}}(A)$ induces a unitary tensor functor $F:\mathcal{C}^{mp}\boxtimes\mathcal{C}\to{\sf Bim_{fgp}}(A^{\text{op}}\otimes A)$, where $\otimes$ is either the minimal or the maximal tensor product of C-algebras. $F$ is constructued from $\tilde{F}$ as follows. First, note that there are equivalences
${\sf Bim_{fgp}}(A^{\text{op}}\underset{\min}{\otimes}A)\simeq{\sf Bim_{fgp}}(A^{\text{op}})\underset{\min}{\boxtimes}{\sf Bim_{fgp}}(A)\ ,\ {\sf Bim_{fgp}}(A^{\text{op}}\underset{\max}{\otimes}A)\simeq{\sf Bim_{fgp}}(A^{\text{op}})\underset{\max}{\boxtimes}{\sf Bim_{fgp}}(A)\ .$
Indeed, letting $\otimes=\underset{\max}{\otimes}$, we have the canonical forgetful functors
${\sf Bim_{fgp}}(A^{\text{op}}\otimes A)\rightarrow\operatorname{Corr}(A^{\text{op}}\otimes A)\simeq\operatorname{Corr}(A^{\text{op}})\boxtimes\operatorname{Corr}(A)\leftarrow{\sf Bim_{fgp}}(A^{\text{op}})\boxtimes{\sf Bim_{fgp}}(A)$
(see [MR4162123], Section 3) and it is easy to see that the essential image of ${\sf Bim_{fgp}}(A^{\text{op}}\otimes A)$ and ${\sf Bim_{fgp}}(A^{\text{op}})\boxtimes{\sf Bim_{fgp}}(A)$ under the corresponding forgetful functors are equivalent.
Given $X\in{\sf Bim_{fgp}}(A)$, let $X^{\#}$ be the complex conjugate vector space. Then
@@
Moreover, if $\psi:(\mathcal{M},m)\to(\mathcal{N},n)$ is a morphism of pointed right $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C-categories, then there is an induced *-homomorphism $\psi_{*}:(A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{m}\to(A^{\text{op}}\otimes A)\underset{F}{\rtimes}\mathcal{Y}_{n}$ that fixes $B_{F}$ point-wise. The operation $\psi\mapsto\psi_{*}$ respects composition, and we obtain therefore a functor from the category of pointed right $\mathcal{Z}\mathrm{Hilb}(\mathcal{C})$-module C-categories to extensions of the symmetric enveloping algebra $B_{F}$.
@@ 6.4. Realization of Categories of Internal Modules
Let $\mathcal{D}$ be a unitary tensor category, $\mathbb{D}$ a C-algebra object in $\operatorname{Vec}(\mathcal{D})$, and suppose $F:\mathcal{D}\to\operatorname{Corr}(B)$ is a fully-faithful unitary tensor functor. Recall that $\mathcal{M}_{\mathbb{D}}$ denotes the Cauchy completion of the category of free left $\mathbb{D}$-modules of the form $\mathbb{D}\odot a=:\mathbb{D}_{a}$, with $a\in\mathbb{D}$.
@@
Given $\xi_{c}\odot\eta_{c}\in F(c)\odot\mathbb{D}(c)$ and $\xi_{b}\odot\eta_{b}\in F(b)\odot\mathbb{D}_{a}(b)$, define
@@
See [Hataishi_Palomares_2025]. Identically to how one describes the algebra structure of $|\mathbb{D}|^{\circ}_{F}$, the following Lemma is proven.
@@
$\langle\xi_{b_{1}}\odot\eta_{b_{1}},\xi_{b_{2}}\odot\eta_{b_{2}}\rangle:=\sum_{c\in\operatorname{Irr}(\mathcal{D})}\sum_{v\in O(c,b_{1}\bar{b}_{2})}F(v^{*})\left(F_{2}(\xi_{b_{1}}\odot j^{F}(\xi_{b_{2}}))\right)\odot\mathbb{D}(v)\left(\mathbb{D}_{2}((\eta_{b_{1}}\odot j^{\mathbb{D}}(\eta_{b_{2}}))(\mathrm{id}_{b_{1}}\odot R_{a}\odot\mathrm{id}_{\bar{b}_{2}}))\right)$
@@
Proof.
All properties but positivity and faithfulness are easy to obtain, argument being identical to the construction of the *-algebra structure of $|\mathbb{D}|_{F}^{\circ}$.
First, observe that for $\xi\odot\eta\in F(b)\odot\mathbb{D}_{a}(b)$ we have $\langle\xi\odot\eta,\xi\odot\eta\rangle\neq 0$, due to the equalities
$\|F_{2}(\xi\odot j^{F}(\xi))\|$ $=\|\xi\underset{B}{\boxtimes}j(\xi)\|>0,$
$\|\mathbb{D}_{2}((\eta\odot j^{\mathbb{D}}(\eta))(\mathrm{id}_{b}\odot R_{a}\odot\mathrm{id}_{\bar{b}}))\|$ $=d_{a}^{1/2}\|\eta\underset{\mathbb{D}({\mathds{1}})}{\boxtimes}j(\eta)\|>0\ .$
The first equality in the second line follows from the fact that $d_{a}^{-1/2}R_{a}$ is an isometry. Moreover, $\langle\xi\odot\eta,\xi\odot\eta\rangle$ is a positive element in $|\mathbb{D}|_{F}^{\circ}$. Indeed,
$F_{2}(\xi\odot j(\xi))\odot\mathbb{D}_{2}(\eta\odot j(\eta))\in F(b\bar{b})\odot\mathbb{D}(b\bar{a}a\bar{b})$
projects to a positive element of $|\mathbb{D}|_{F}^{\circ}$. Now, $\mathbb{D}(\mathrm{id}_{b}\odot R_{a}^{*}\odot\mathrm{id}_{\bar{b}}):\mathbb{D}(b\bar{a}a\bar{b})\to\mathbb{D}(b\bar{b})$ is a completely positive map, and therefore the projection of
$F_{2}(\xi\odot j(\xi))\odot\mathbb{D}_{2}(\eta\odot j(\eta)(\mathrm{id}_{b}\odot R_{a}\odot\mathrm{id}_{\bar{b}}))=(\mathrm{id}\odot\mathbb{D}(\mathrm{id}_{b}\odot R_{a}^{*}\odot\mathrm{id}_{\bar{b}}))\left(F_{2}(\xi\odot j(\xi))\odot\mathbb{D}_{2}(\eta\odot j(\eta))\right)$
is also positive in $|\mathbb{D}|_{F}^{\circ}$.
From the additivity of both $F$ and $\mathbb{D}$, the claimed properties of $\langle\cdot,\cdot\rangle$ follow from the above by an application of a standard amplification trick: seeing vectors in $\oplus_{b}F(b)\odot\mathbb{D}_{a}(b)$ as vectors in $F(X)\odot\mathbb{D}_{a}(X)$, where
$X=\bigoplus_{i\in I}b_{i}$
for a finite indexing set $I$. ∎
@@
Proof.
Indeed, the right given $A$-action is a direct sum of amplifications of the original right $A$ actions on the modules $F(b)$. ∎
The following Lemma can be verified straightforwardly.
@@
in $B\otimes\mathbb{D}({\mathds{1}})$ by faithfulness of the canonical conditional expectation. There exists a state $\omega$ on $\mathbb{D}({\mathds{1}})$ such that
@@
Then $\Im(\tilde{\iota}_{\omega})\subset\Im(\iota_{\omega})$. The inequality $(\mathrm{id}_{B}\otimes\omega)(\mathbb{E}(\langle T(\zeta),T(\zeta)\rangle))>0$ now implies that $\tilde{\iota}_{\omega}(T)\neq 0$, so $T$ is a non-zero operator in $\Im(\iota_{\omega})\simeq\mathcal{M}_{\mathbb{D}}(\mathbb{D}_{a_{1}},\mathbb{D}_{a_{2}})$. ∎